### v1

# Invariant Subspaces of Perturbed Backward Shift

Abstract.
We represent closed subspaces of the Hardy space that are invariant under finite-rank perturbations of the backward shift. We apply this to classify almost invariant subspaces of the backward shift and represent closed subspaces that are invariant under a more refined version of nearly invariant subspaces of the backward shift. Kernels of certain perturbed Toeplitz operators are examples of the newly introduced nearly invariant subspaces.
Key words and phrases:
Hardy space, invariant subspace, shift operator, nearly invariant subspace, model space, Schmidt subspace
2010 Mathematics Subject Classification:
30H10, 47A15, 47B35, 46E22
1. Introduction
In this paper, we bring together three problems concerning invariant subspaces. These problems have one commonality: the backward shift operator $T_{z}^{*}$ acting on $H^{2}$ appears in all of them. Recall the standard notation that $T_{z}$ is the multiplication operator by the coordinate function $z$; alternatively, the Toeplitz operator with the symbol $z$; and $H^{2}$ is the Hardy space of all square summable analytic functions on the open unit disc $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$ (see (2.3) for more details). Therefore
$T_{z}f=zf,$
for all $f\in H^{2}$. We effectively connect invariant subspaces of three types of operators: finite rank perturbations of $T_{z}^{*}$, nearly invariant subspaces of $T_{z}^{*}$ (in a larger sense), and almost invariant subspaces of $T_{z}^{*}$. Strategically, we first represent closed invariant subspaces of finite rank perturbations of $T_{z}^{*}$, and then we use that technique to represent and classify the invariant subspaces of the two subsequent classes of operators.
Therefore, the broad topic of this paper is invariant subspaces of $T_{z}^{*}$-related operators. Linear analysis has been contending with invariant subspaces from the very beginning. The famous “invariant subspace problem” calls for finding a nontrivial closed subspace that is invariant under a bounded linear operator on a Hilbert space. This problem is still open. Even worse, the existence of an invariant subspace for rank-one perturbations of diagonal operators is unknown [5]. In the same vein, the lattice of the invariant subspaces of simple composition operators is regarded as a difficult problem and is inextricably linked to the original invariant subspace problem [11]. Hence, identifying the lattices of invariant subspaces of natural operators is commonly a complex problem, and its solution often provides a valuable understanding of various topics (cf. Beurling theorem [1] and it’s applications [12, 13]). In this paper, we will also address all of these major issues related to $T_{z}^{*}$.
Before moving on, it is worth remarking that the results and approach of this paper are an enhancement of the circles of ideas that were presented by Hitt [9], which were further fine-tuned by Sarason. In fact, we identify Sarason’s observation on a specific rank-one projection [15, page 482] (also see (1.2) for more details) and use that perspective to provide full-length parametrizations of invariant subspaces of all finite-rank perturbations of $T^{*}_{z}$. We have subsequently laid out the road map to link the other two ideas.
To follow the progression of ideas, we begin with perturbations of $T_{z}^{*}$. Each pair of nonzero vectors $u$ and $v$ in $H^{2}$ yields the rank-one operator $u\otimes v$, where
$(u\otimes v)f=\langle f,v\rangle u,$
for all $f\in H^{2}$. Similarly, rank-$m$ operators are those whose ranges are $m$-dimensional. They can be represented as:
$\sum_{i=1}^{m}v_{i}\otimes u_{i},$
where $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets of vectors in $H^{2}$. For the remainder of this section, we fix these orthonormal sets. We are interested in the invariant subspaces of a general finite-rank perturbation $X$ of $T_{z}^{*}$, where
$X=T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}.$
In Corollary 2.2, we present a complete description of invariant subspaces of $X$:
Theorem 1.1.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ such that
$\mathcal{M}=[F_{0},1]_{1\times(p+1)}\mathcal{K}_{\Psi},$
where $\mathcal{K}_{\Psi}=H^{2}_{\mathbb{C}^{p+1}}/\Psi H^{2}_{\mathbb{C}^{p^{\prime}}}$ is the model space, $F_{0}=[f_{1},\ldots,f_{p}]$, and $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|f_{0}\|^{2}=\|F_{0}F+f_{0}\|^{2}\}.$
In the given context, $H^{2}_{\mathbb{C}^{p}}$ represents the $\mathbb{C}^{p}$-valued Hardy space over $\mathbb{D}$, which may also be seen as $p$-copies of $H^{2}$. Given Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$, we denote $\mathcal{B}(\mathcal{H},\mathcal{K})$ as the space of all bounded linear operators (and write $\mathcal{B}(\mathcal{H})$ if $\mathcal{K}=\mathcal{H}$). The symbol $H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ denotes the set of $\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})$-valued bounded analytic functions on $\mathbb{D}$, where a function $\Phi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ is inner if
$\Phi(z)^{*}\Phi(z)=I_{\mathbb{C}^{p^{\prime}}}\qquad(z\in\mathbb{T}\;a.e.)$
Moreover, $[F_{0},1]_{1\times(p+1)}=[f_{1},\ldots,f_{p},1]$ is a row vector, and $\mathcal{K}_{\Psi}$’s elements are columns.
In fact, we prove Theorem 1.1 in a much higher generality (see Theorem 2.1): we simply replace the backward shift $T_{z}^{*}$ by the adjoint of an analytic Toeplitz operator $T_{\varphi}$, where the symbol $\varphi\in H^{\infty}$ is an inner function vanishing at the origin.
Before proceeding, we need the definition of nearly $T_{z}^{*}$-invariant subspaces (see Hayashi [8], Hitt [9], and Sarason [15]). A closed subspace $\mathcal{M}$ of $H^{2}$ is nearly $T_{z}^{*}$-invariant if
$T_{z}^{*}f\in\mathcal{M},$
for all $f\in\mathcal{M}\cap zH^{2}$. Now, we will provide a few comments regarding Theorem 1.1. First, this result deals with linear perturbations, a theory in and of itself that is well-known to be challenging (see the classic [10]). Second, this result parameterizes all of the invariant subspaces of finite rank perturbations of $T_{z}^{*}$, which was known to be a challenging goal. As far as we know, this is the first result in this direction. Third, we provide this result in the tradition of Hitt and Sarason’s classical argumentation. Indeed, Sarason showed in [15, page 481] that if $\mathcal{M}$ is a nearly $T_{z}^{*}$-invariant subspace then there is a unique function
$g\in\mathcal{M}\ominus(\mathcal{M}\cap zH^{2}),$ (1.1)
such that $\mathcal{M}$ is invariant under rank-one perturbation $T_{z}^{*}-T_{z}^{*}g\otimes g$, that is
$(T_{z}^{*}-T_{z}^{*}g\otimes g)\mathcal{M}\subseteq\mathcal{M}.$ (1.2)
Sarason’s observation motivated us to attempt full-length representations of invariant subspaces of finite-rank perturbations, leading to the formation of Theorem 1.1. At the very start of this section, this point was briefly mentioned. Finally, we recall the representation of nearly $T_{z}^{*}$-invariant subspaces, which is attributed to Hitt [9] (also see Sarason [15]). Every nearly $T_{z}^{*}$-invariant subspace $\mathcal{M}\subseteq H^{2}$ admits the representation
$\mathcal{M}=gH^{2},$
where $g$ is a unique function as in (1.1). Undoubtedly, the result in Theorem 1.1 is in line with the same fundamentals. As previously said, in the present case, we have effectively applied the core ideas that were first put forward by Hitt and Sarason a few decades ago.
Now we turn to almost invariant subspaces for $T_{z}^{*}$. In the setting of the backward shift operator, this was introduced by Chalendar, Gallardo-Gutiérrez, and Partington [4]. Let $T\in\mathcal{B}(\mathcal{H})$, and let $\mathcal{M}\subseteq\mathcal{H}$ be a closed subspace. We say that $\mathcal{M}$ is almost invariant for $T$ if there exists a finite-dimensional subspace $\mathcal{F}\subseteq\mathcal{H}$ such that
$T\mathcal{M}\subseteq\mathcal{M}\oplus\mathcal{F}.$
The defect of the space $\mathcal{M}$ is the lowest possible dimension of such a space $\mathcal{F}$. Almost invariant subspaces for $T_{z}^{*}$ were classified in [4, Corollary 3.4] in terms of nearly invariant subspaces with finite defect indices. We also study the same thing, but we take a completely different approach: we enter into this problem through finite-rank perturbations of $T_{z}^{*}$. Indeed, a special version of Theorem 5.1 implies the following link: almost invariant subspaces and invariant subspaces of finite rank perturbations of the backward shift operator.
Theorem 1.2.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$, and let $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ be orthonormal sets in $H^{2}$. If $\mathcal{M}$ is invariant under
$T_{z}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i},$
then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect dimension at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect dimension $m$, then $\mathcal{M}$ is invariant under
$T_{z}^{*}-\sum_{i=1}^{m}f_{i}\otimes T_{z}f_{i},$
where $\{f_{i}:1\leq i\leq m\}$ is an orthonormal basis of the $m$ dimensional defect space.
This yields a complete characterization of almost invariant subspaces of $T_{z}^{*}$ (see Corollary 5.2):
Theorem 1.3.
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. If $\mathcal{M}$ is almost invariant under $T_{z}^{*}$, then there exist natural numbers $p^{\prime}<p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
$\mathcal{M}=[F_{0},1]_{1\times(p+1)}\mathcal{K}_{\Psi},$ (1.3)
where $F_{0}=[f_{1},\ldots,f_{p}]$, and $\{f_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}(u_{i}):i=1,\ldots,m\}$, and $\{u_{i}\}_{i=1}^{m}$ is an orthonormal basis for the defect space. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|{F}\|^{2}+\|{f_{0}}\|^{2}=\|F_{0}F+f_{0}\|^{2}\},$ (1.4)
and
$\mathcal{J}(k_{1},\ldots,k_{p+1})=f_{1}k_{1}+\cdots+f_{p}k_{p}+k_{p+1},$ (1.5)
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$, defines a unitary operator $\mathcal{J}:\mathcal{K}_{\Psi}\to\mathcal{M}$. Conversely, if $\mathcal{M}$ has the representation (1.3) for some orthonormal basis $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ with $\mathcal{K}_{\Psi}$ as in (1.4), and a unitary $\mathcal{J}$ as in (1.5), then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$.
It should be noted that Chalendar, Gallardo-Gutiérrez, and Partington were the first to classify almost invariant subspaces under $T_{z}^{*}$ [4, Corollary 3.4]. However, the above result differs significantly from [4], which compares nearly $T_{z}^{*}$-invariant subspaces of finite defect indices with almost invariant subspaces. Here, we connect them with finite rank perturbations of $T_{z}^{*}$. This, in particular, yields more concrete representations of the backward shift-invariant subspace $\mathcal{K}_{\Psi}$. See the end of Section 5 for more comments.
Now we turn to a more refined version of nearly $T_{z}^{*}$-invariant subspaces. The motivation for this arises partly from the matching theory of finite-rank perturbations of $T_{\varphi}^{*}$ as well as the potential of the theory of kernels of finite-rank perturbations of Toeplitz operators, which we will elaborate on shortly. We start by setting up the notations. Let $\{z_{k}\}_{k=1}^{n}$ be a sequence of points (possibly repeated) in $\mathbb{D}$, and assume that
$z_{j}=0,$
for some $j=1,\ldots,n$. Define the corresponding Blaschke product $B_{n}$ by
$B_{n}(z)=\prod_{k=1}^{n}\dfrac{z_{k}-z}{1-\bar{z}_{k}z},$
for all $z\in\mathbb{D}$. Clearly, $B_{n}\in H^{\infty}$ is an inner function, and
$B_{n}(0)=0.$
We will consider this Blaschke product for the rest of the discussion in this section.
Definition 1.4.
A closed subspace $\mathcal{M}\subseteq H^{2}$ is said to be nearly $T^{*}_{z,B_{n}}$-invariant if
$T_{z}^{*}f\in\mathcal{M},$
for all
$f\in\mathcal{M}\cap B_{n}H^{2}.$
Evidently, nearly $T^{*}_{z,B_{1}}$-invariant subspaces are precisely nearly $T_{z}^{*}$-invariant subspaces. In Theorem 6.3, we present a complete description of nearly $T^{*}_{z,B_{n}}$-invariant subspaces of $H^{2}$:
Theorem 1.5.
Let $\mathcal{M}$ be a nontrivial closed subspace of $H^{2}$. Then $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant if and only if there exist natural numbers $r^{\prime}\leq r\leq n$ and an inner function $\Theta\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{r^{\prime}},\mathbb{C}^{r})}$ such that $\Theta(0)=0$ and
$\mathcal{M}=[g_{1},\ldots,g_{r}]_{1\times r}\mathcal{K}_{\Theta},$
where $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B{H}^{2})$, and
$\mathcal{J}(k_{1},\ldots,k_{r})=g_{1}k_{1}+\cdots+g_{r}k_{r},$
for all $(k_{1},\ldots,k_{r})\in\mathcal{K}_{\Theta}$, defines a unitary operator $\mathcal{J}:\mathcal{K}_{\Theta}\rightarrow\mathcal{M}$.
Kernels of Toeplitz operators are crucial in the context of examples of nearly $T_{z}^{*}$-invariant subspaces. In a comparable manner, we provide examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. Curiously, certain dedicated finite-rank perturbations of Toeplitz operators result in nontrivial examples of $T^{*}_{z,B_{n}}$-invariant subspaces. More specifically, given the symbol $\varphi\in L^{\infty}$ (the space of all essentially bounded functions on $\mathbb{T}$), we follow the standard notation of the Toeplitz operator as $T_{\varphi}$, where
$T_{\varphi}f=P_{H^{2}}(\varphi f),$
for all $f\in H^{2}$, and $P_{H^{2}}$ denotes the orthogonal projection (or, the Szegö projection) from $L^{2}$ (the space of all square-integrable functions on $\mathbb{T}$) onto $H^{2}$. Pick an arbitrary orthonormal basis $\{f_{i}\}_{i=1}^{n}$ of $\mathcal{K}_{B_{n}}$, and consider the finite-rank perturbation of $T_{\varphi}$ as $T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i}$. In Theorem 7.1, we prove that
$\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i}),$
is nearly $T^{*}_{z,B_{n}}$-invariant subspace. We are of the opinion that this space is significant. We refer to [3, 7] for information regarding the kernels of Toeplitz operators and nearly $T_{z}^{*}$-invariant subspaces.
We also consider representations of invariant subspaces of finite-rank perturbations of $T_{z}$. We recall a general fact from elementary functional analysis. Let $\mathcal{S}$ be a closed subspace of a Hilbert space $\mathcal{H}$ and let $T\in\mathcal{B}(\mathcal{H})$. Then $\mathcal{S}$ is invariant under $T$ if and only if $\mathcal{S}^{\perp}$ is invariant under $T^{*}$. In particular, representations of $T$-invariant subspaces also yield representations of $T^{*}$-invariant subspaces. Nevertheless, the explicit feature of the invariant property (if any) may not be maintained when representations are transmitted from one to another, particularly in a concrete situation such as ours. Here, we are talking about the representation of invariant subspaces of finite-rank perturbations of the forward shift $T_{z}$, as we are already aware of the same for finite-rank perturbations of $T_{z}^{*}$.
We identify a class of perturbations of $T_{z}$ for which one can say more concretely about the lattice of invariant subspaces (see Theorem 4.1):
Theorem 1.6.
Let $\mathcal{M}$ be an invariant subspace of $T_{z}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$. Suppose
$\mathcal{L}:=span\{P_{\mathcal{M}^{\perp}}v_{1},\ldots,P_{\mathcal{M}^{\perp}}v_{m}\},$
and let $p=\dim\mathcal{L}$. If $\{\varphi_{i}\}_{i=1}^{p}\subseteq H^{\infty}$ is an orthonormal basis for $\mathcal{L}$, then there exists an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ for some $p^{\prime}\leq p+1$ such that
=$\mathcal{M}=\{g\in H^{2}:(T_{\bar{\varphi}_{1}}g,\ldots,T_{\bar{\varphi}_{p}}g,g)\in\Psi H^{2}_{\mathbb{C}^{p^{\prime}}}\}.$
Some of the results stated above are merely one-directional. This is also in keeping with Hitt and Sarason’s perspectives on representations of nearly $T_{z}^{*}$-invariant subspaces. In some of the instances, we present a converse directional result that is restricted to a particular class of finite-rank perturbations. These perturbations are motivated by Sarason’s rank-one perturbation, as pointed out in (1.2). and typically they look like
$T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}f_{i}\otimes f_{i}.$ (1.6)
We call them Sarason-type perturbations.
The remaining sections of the paper are structured in the following manner: Section 2 is the heart of the paper, which gives invariant subspaces of finite rank perturbations of Toeplitz operators with inner symbols vanishing at 00. In Section 3, we apply the results of the preceding section to the particular situation of the forward shift operator. We additionally take into account the Sarason-type perturbations. Section 4 deals with invariant subspaces of certain finite-rank perturbations of $T_{z}$. In Section 5, we continue our approach to finite-rank perturbations and determine the lattice of almost invariant subspaces of $T_{z}^{*}$. Section 6 introduces the concept of nearly $T^{*}_{z,B_{n}}$-invariant subspaces and presents an invariant subspace theorem. The final section, Section 7, is devoted to examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. These subspaces are precisely the kernels of certain finite-rank perturbations of Toeplitz operators.
2. Perturbations of $T_{\varphi}^{*}$
In this section, we aim to represent invariant subspaces of finite-rank perturbations of $T_{z}^{*}$. However, our techniques are more refined, allowing us to produce results with a significantly higher level of generality. In the following, we establish the problem within the context of the Toeplitz operator with an inner function symbol that vanishes at the origin. Recall that a function $\varphi\in H^{\infty}$ is inner if
$|\varphi(z)|=1,$
for all $z\in\mathbb{T}$ a.e. (in the sense of radial limits of $H^{\infty}$-functions). The analytic Teoplitz operator with symbol $\varphi\in H^{\infty}$ is then defined by
$T_{\varphi}f=\varphi f,$
for all $f\in H^{2}$. We will consider rank-$m$ perturbations of the Toepliz operator $T_{\varphi}$. Let $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ be arbitrary but fixed orthonormal sets of vectors in $H^{2}$. Define the rank-$m$ perturbation $\mathcal{T}_{[\varphi;u,v]}$ of $T_{\varphi}^{*}$ on $H^{2}$ by
$\mathcal{T}_{[\varphi;u,v]}=T_{\varphi}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}.$
Our goal is to parameterize all the invariant subspaces of $\mathcal{T}_{[\varphi;u,v]}$.
To achieve this, we first recall that an operator $T\in\mathcal{B}(\mathcal{H})$ is said to be in $C_{\cdot 0}$ (in short, $T\in C_{\cdot 0}$) if $T$ is a contraction and
$\text{SOT}-\lim_{n\rightarrow\infty}T^{*n}=0,$
that is, ${T^{*}}^{n}h\to 0$ as $n\to\infty$ for all $h\in\mathcal{H}$. The following fact is particularly useful, as it affirms that $C_{\cdot 0}$ is invariant under multiplication by projections with finite-codimenional ranges (see [2, Lemma 3.3]): Let $T$ on $\mathcal{H}$ be a $C_{\cdot 0}$ contraction, and let $\mathcal{F}$ be a finite-codimensional subspace of $\mathcal{H}$. If $\dim(\text{ran}(I-T^{*}T))<\infty$, then
$TP_{\mathcal{F}}\in C_{\cdot 0}.$ (2.1)
Here, given a closed subspace $\mathcal{S}$ of $\mathcal{H}$, we denote by $P_{\mathcal{S}}$ the orthogonal projection onto $\mathcal{S}$.
Returning to our perturbation situation, we first define Sarason-type finite-rank perturbation $\mathcal{T}_{[\varphi,u]}$ on $H^{2}$ as
$\mathcal{T}_{[\varphi,u]}:=T_{\varphi}^{*}-\sum_{i=1}^{m}T_{\varphi}^{*}u_{i}\otimes u_{i}.$
In other words, $\mathcal{T}_{[\varphi,u]}=T_{\varphi}^{*}(I-\sum_{i=1}^{m}u_{i}\otimes u_{i})$. Equivalently
$\mathcal{T}_{[\varphi,u]}f=T_{\varphi}^{*}f-\sum_{i=1}^{m}\langle f,u_{i}\rangle T_{\varphi}^{*}u_{i},$
for all $f\in H^{2}$. We claim that
$\mathcal{T}_{[\varphi,u]}^{*}\in C_{\cdot 0}.$ (2.2)
This is essentially the result of the fact (2.1). Indeed, first we observe that
$\mathcal{T}_{[\varphi,u]}=T_{\varphi}^{*}(I-P_{\mathcal{F}}),$
where $\mathcal{F}=span\{u_{1},u_{2},\ldots,u_{m}\}$ is a finite-dimensional subspace of $H^{2}$. Since $T_{\varphi}$ is a contraction and $I-P_{\mathcal{F}}$ is an orthogonal projection, the operator $\mathcal{T}_{[\varphi,u]}$ is a contraction. Since $T_{\varphi}\in C_{\cdot 0}$ and $\mathcal{F}$ is finite-dimensional, (2.1) stated above implies that $T_{\varphi}(I-P_{\mathcal{F}})$ is a $C_{\cdot 0}$ contraction. For each $g\in H^{2}$ and $n\geq 1$, we compute
$\mathcal{T}_{[\varphi,u]}^{n}g=(T_{\varphi}^{*}(I-P_{\mathcal{F}}))^{n}g=T_{\varphi}^{*}([T_{\varphi}(I-P_{\mathcal{F}})]^{*})^{n-1}(I-P_{\mathcal{F}})g.$
Then
$\|\mathcal{T}_{[\varphi,u]}^{n}g\|\leq\|T_{\varphi}^{*}\|\|([T_{\varphi}(I-P_{\mathcal{F}})]^{*})^{n-1}(I-P_{\mathcal{F}})g\|,$
implies that $\mathcal{T}_{[\varphi,u]}^{*}\in C_{\cdot 0}$, thereby proving (2.2). We emphasize that the technique above has become typical in showing similar results.
We need to fix some more notations. Given a natural number $p\geq 1$, $H^{2}_{\mathbb{C}^{p}}$ will denote the $\mathbb{C}^{p}$-valued Hardy space over $\mathbb{D}$. It is defined as:
$H^{2}_{\mathbb{C}^{p}}=\Big{\{}F(z)=\sum_{n=0}^{\infty}A_{n}z^{n}:\|F\|^{2}=\sum_{n=0}^{\infty}\|{A_{n}}\|_{\mathbb{C}^{p}}^{2}<\infty,A_{n}\in\mathbb{C}^{p},z\in\mathbb{D}\Big{\}}.$ (2.3)
Furthermore, in the subsequent discussion, we shall define identification as up to unitary equivalence. We will frequently identify $H^{2}_{\mathbb{C}^{p}}$ with $H^{2}\otimes\mathbb{C}^{p}$. The other common identification that will be used throughout is $H^{2}_{\mathbb{C}^{p+1}}$ and $H^{2}_{\mathbb{C}^{p}}\oplus H^{2}$. The concept of model spaces is an additional component that will be utilized in the contents that follow. Given $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$, the quotient space
$\mathcal{K}_{\Psi}=H^{2}_{\mathbb{C}^{p}}/\Psi H^{2}_{\mathbb{C}^{p^{\prime}}},$ (2.4)
classifies $(T_{z}^{*}\otimes I_{\mathbb{C}^{p}})$-invariant closed subspaces of $H^{2}_{\mathbb{C}^{p}}$. This is the consequence of the celebrated Beurling-Lax theorem. The subspace $\mathcal{K}_{\Psi}$ is commonly referred to as the model space corresponding to the inner function $\Psi$.
Now we are in a position to present the description of invariant subspaces of $T_{\varphi}^{*}$. In the following statement, $[f_{1},\ldots,f_{p},1]_{1\times(p+1)}$ refers to the row vector in the $(p+1)$-copies of $H^{2}$ (which is identified with $H^{2}_{\mathbb{C}^{p+1}}$). Moreover, often we will express vectors $F$ in $H^{2}_{\mathbb{C}^{p}}$ as column vectors $[f_{1},\ldots,f_{p}]^{t}$, where $f_{i}\in H^{2}$ for all $i=1,\ldots,p$.
Theorem 2.1.
Let $\varphi\in H^{\infty}$ be an inner function, and let $\mathcal{M}$ be a closed subspace of $H^{2}$. Suppose $\varphi(0)=0$. If $\mathcal{M}\subseteq H^{2}$ is an invariant subspace of $T_{\varphi}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exists a $(T_{\varphi}^{*}\otimes I_{\mathbb{C}^{p+1}})$-invariant closed subspace $\mathcal{K}\subseteq H^{2}_{\mathbb{C}^{p+1}}$ such that
$\mathcal{M}=[F_{0},1]_{1\times(p+1)}\mathcal{K},$
where $F_{0}=[f_{1},\ldots,f_{p}]_{1\times p}$, and $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for
$\text{span}\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}.$
Moreover, $\mathcal{K}$ can be represented as
$\mathcal{K}:=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|f_{0}\|^{2}=\|(F,f_{0})\|^{2}\}.$
Proof.
We open by noting that the following inequality occurs within the context of the theorem’s statement:
$p:=\dim\mathcal{W}\leq m=\text{rank }(\sum_{i=1}^{m}v_{i}\otimes u_{i}),$
where following our convention, we simply write $\mathcal{T}_{[\varphi;u,v]}=T_{\varphi}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, and define the subspace $\mathcal{W}:=span\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Assume that $\mathcal{M}$ is invariant under $\mathcal{T}_{[\varphi;u,v]}$. If $\mathcal{W}=\{0\}$, then $\{u_{1},u_{2},\ldots,u_{m}\}\perp\mathcal{M}$, which implies that $\mathcal{M}$ is $T_{\varphi}^{*}$ invariant. In this case, we simply write $\mathcal{M}=[1]\mathcal{K}$, where $\mathcal{K}=\mathcal{M}$. Next, we assume that
$\mathcal{W}\neq\{0\}.$
Let
$dim(\mathcal{W})=p.$
Then $1\leq p\leq m$. Pick an orthonormal basis $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, and denote the row operator
$F_{0}=[f_{1},f_{2},\ldots,f_{p}]_{1\times p}.$
Decompose $\mathcal{M}$ as
$\mathcal{M}=\mathcal{W}\oplus(\mathcal{M}\cap\mathcal{W}^{\perp}).$
Fix $f\in\mathcal{M}$ and write
$f=P_{\mathcal{W}}(f)\oplus P_{\mathcal{M}\cap\mathcal{W}^{\perp}}(f)=(a_{01}f_{1}+\cdots+a_{0p}f_{p})\oplus g_{1},$
that is
$f=F_{0}A_{0}\oplus g_{1},$ (2.5)
where
$A_{0}=[a_{01},\ldots,a_{0p}]^{t}\in\mathbb{C}^{p},$
and
$g_{1}:=P_{\mathcal{M}\cap\mathcal{W}^{\perp}}f\in\mathcal{M}\cap\mathcal{W}^{\perp}.$
Moreover
$\|{f}\|^{2}=\|{A_{0}}\|^{2}+\|{g_{1}}\|^{2}.$
Now we proceed to simplify the function $g_{1}$. As $\mathcal{T}_{[\varphi;u,v]}\mathcal{M}\subseteq\mathcal{M}$, it follows that
$l_{1}:=\mathcal{T}_{[\varphi;u,v]}g_{1}\in\mathcal{M}.$
We also compute
$l_{1}=(T_{\varphi}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i})g_{1}=T_{\varphi}^{*}g_{1}-\sum_{i=1}^{m}\langle g_{1},u_{i}\rangle v_{i}=T_{\varphi}^{*}g_{1},$
as $g_{1}\in\mathcal{M}\cap\mathcal{W}^{\perp}$ and $\mathcal{W}=\text{span}\{P_{\mathcal{M}}u_{j}:j=1,\ldots,m\}$. Since $T_{\varphi}T_{\varphi}^{*}=I-P_{\mathcal{K}_{\varphi}}$, we conclude that $T_{\varphi}l_{1}=T_{\varphi}T_{\varphi}^{*}g_{1}=g_{1}-P_{\mathcal{K}_{\varphi}}g_{1}$. This readily implies that
$g_{1}=\varphi l_{1}+P_{\mathcal{K}_{\varphi}}g_{1},$
and then $f=F_{0}A_{0}\oplus(l_{1}\varphi+P_{\mathcal{K}_{\varphi}}g_{1})$. Also, we have
$\|{g_{1}}\|^{2}=\|{l_{1}}\|^{2}+\|P_{\mathcal{K}_{\varphi}}g_{1}\|^{2}.$
By summarizing all the above observations, we conclude that
$\begin{cases}f=F_{0}A_{0}+l_{1}\varphi+P_{\mathcal{K}_{\varphi}}g_{1}\\ \|{f}\|^{2}=\|{A_{0}}\|^{2}+\|{l_{1}}\|^{2}+\|P_{\mathcal{K}_{\varphi}}g_{1}\|^{2},\end{cases}$
and
$l_{1}=(\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}})f.$
We now employ the technique of representing $f\in\mathcal{M}$ outlined above to represent $l_{1}\in\mathcal{M}$ (which will lead to the induction steps). Therefore, we first decompose
$l_{1}=F_{0}A_{1}\oplus g_{2},$
where $g_{2}=P_{\mathcal{M}\cap\mathcal{W}^{\perp}}l_{1}\in\mathcal{M}\cap\mathcal{W}^{\perp}$, $A_{1}\in\mathbb{C}^{p}$, and
$\|{l_{1}}\|^{2}=\|{A_{1}}\|^{2}+\|{g_{2}}\|^{2}.$
A computation similar to the one above implies that
$g_{2}=l_{2}\varphi+P_{\mathcal{K}_{\varphi}}g_{2},$
and
$\|{g_{2}}\|^{2}=\|{l_{2}}\|^{2}+\|P_{\mathcal{K}_{\varphi}}g_{2}\|^{2},$
where
$l_{2}:=\mathcal{T}_{[\varphi;u,v]}g_{2}\in\mathcal{M}.$
Therefore, after these two steps, we have
$\begin{cases}f=F_{0}(A_{0}+A_{1}\varphi)+l_{2}\varphi^{2}+(P_{\mathcal{K}_{\varphi}}g_{1}+(P_{\mathcal{K}_{\varphi}}g_{2})\varphi)\\ \|{f}\|^{2}=\|{A_{0}}\|^{2}+\|{A_{1}}\|^{2}+\|{l_{2}}\|^{2}+\|P_{\mathcal{K}_{\varphi}}g_{1}\|^{2}+\|P_{\mathcal{K}_{\varphi}}g_{2}\|^{2}.\end{cases}$
We also have $l_{2}=\mathcal{T}_{[\varphi;u,v]}g_{2}=\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}l_{1}$, and then
$l_{2}=\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}\mathcal{T}_{[\varphi;u,v]}g_{1}=\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}f,$
that is
$l_{2}=(\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}})^{2}f.$
By continuing this process, we obtain
$\begin{cases}f=F_{0}(\sum_{j=0}^{n}A_{j}z^{j})+l_{n+1}\varphi^{n+1}+\sum_{j=1}^{n+1}(P_{\mathcal{K}_{\varphi}}g_{j})\varphi^{j-1}\\ \|{f}\|^{2}=\sum_{i=0}^{n}\|{A_{i}}\|^{2}+\|{l_{n+1}}\|^{2}+\sum_{j=1}^{n+1}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2},\end{cases}$ (2.6)
and
$l_{n}=(\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}})^{n}f,$
and $\{A_{i}\}_{i=1}^{n}\subseteq\mathbb{C}^{r}$ for all $n\geq 1$. However, since $\mathcal{T}_{[\varphi;u,v]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}=\mathcal{T}_{[\varphi,u]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}}$, it follows that
$l_{n}=(\mathcal{T}_{[\varphi,u]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}})^{n}f,$
for all $n\geq 1$. By (2.2), we know, in particular, that $\mathcal{T}_{[\varphi,u]}^{*}\in C_{\cdot 0}$ and hence (2.1) implies
$\mathcal{T}_{[\varphi,u]}^{*}(I-P_{\mathcal{W}})\in C_{\cdot 0}.$
Also, it is clear that $P_{\mathcal{M}\cap\mathcal{W}^{\perp}}h=(I-P_{\mathcal{W}})h$ for all $h\in\mathcal{M}$. As
$\|{l_{n+1}}\|=\|{(\mathcal{T}_{[\varphi,u]}P_{\mathcal{M}\cap\mathcal{W}^{\perp}})^{m+1}f}\|\leq\|{\mathcal{T}_{[\varphi,u]}}\|\|{(P_{\mathcal{M}\cap\mathcal{W}^{\perp}}\mathcal{T}_{[\varphi,u]})^{m}f}\|,$
for all $n\geq 1$, we conclude that $\|{l_{n+1}}\|\to 0$ as $n\to\infty$. By (2.6), we know that $\|{f}\|^{2}=\sum_{i=0}^{n}\|{A_{i}}\|^{2}+\|{l_{n+1}}\|^{2}+\sum_{j=1}^{n+1}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2}$, and hence
$\max\Big{\{}\sum_{i=0}^{n}\|{A_{i}}\|^{2},\sum_{j=1}^{n+1}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2}\Big{\}}\leq\|{f}\|^{2},$
yielding
$\max\Big{\{}\sum_{i=0}^{\infty}\|{A_{i}}\|^{2},\sum_{j=1}^{\infty}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2}\Big{\}}\leq\|{f}\|^{2},$
Define
$F(z)=\sum_{i=0}^{\infty}A_{i}\varphi^{i},$
and
$f_{0}(z)=\sum_{j=1}^{\infty}(P_{\mathcal{K}_{\varphi}}g_{j})\varphi^{j-1},$
for all $z\in\mathbb{D}$. We claim that $F\in H^{2}_{\mathbb{C}^{p}}$ and $f_{0}\in H^{2}$. Observe that $\ker T_{\varphi}^{*}=\mathcal{K}_{\varphi}$. Because $T_{\varphi}$ is a pure isometry (that is, a shift operator), by the von Neumann-Wold decomposition theorem, we have the orthogonal decomposition
$H^{2}=\mathcal{K}_{\varphi}\oplus\varphi\mathcal{K}_{\varphi}\oplus\varphi^{2}\mathcal{K}_{\varphi}\oplus\cdots.$
Since $\sum_{j=0}^{\infty}\|P_{\mathcal{K}_{\varphi}}g_{j}\|^{2}<\infty$, in view of the above decomposition, we conclude that $f_{0}\in H^{2}$. For $F$, we set
$T_{\Phi}=T_{\varphi}\otimes I_{\mathbb{C}^{p}}.$
Since $\varphi(0)=0$ implies $\mathbb{C}\subseteq\mathcal{K}_{\varphi}$, we infer that
$\mathbb{C}^{p}\cong\mathbb{C}\otimes\mathbb{C}^{p}\subseteq\mathcal{K}_{\Phi},$
where $\mathcal{K}_{\Phi}=\ker T_{\Phi}^{*}$. Again, since $T_{\Phi}$ on $H^{2}\otimes\mathbb{C}^{p}$ is a pure isometry (as $T_{\Phi}^{*m}=T^{*m}_{\varphi}\otimes I_{\mathbb{C}^{p}}\rightarrow 0$ in SOT), we again have the von Neumann-Wold decomposition as
$H^{2}\otimes\mathbb{C}^{p}=\mathcal{K}_{\Phi}\oplus\Phi\mathcal{K}_{\Phi}\oplus\Phi^{2}\mathcal{K}_{\Phi}\oplus\cdots.$
Because $\mathbb{C}^{p}\subseteq\mathcal{K}_{\Phi}$ and $\sum_{i=0}^{\infty}\|{A_{i}}\|^{2}<\infty$, we conclude that $F\in H^{2}_{\mathbb{C}^{p}}$. Next, for each $n\geq 1$, we consider the $n$-th partial sums
$K_{n}=\sum_{i=0}^{n}A_{i}\varphi^{i},$
and
$k_{n}=\sum_{i=0}^{n}P_{K_{\varphi}}(g_{i+1})\varphi^{i}.$
The representation of $f$ in (2.6) then implies that
$\|{f-F_{0}K_{n}-k_{n}}\|_{2}=\|{l_{n+1}}\|_{2}\to 0,$
in $H^{2}$, and hence
$\|{f-F_{0}K_{n}-k_{n}}\|_{1}\leq\|{f-F_{0}K_{n}-k_{n}}\|_{2}\rightarrow 0,$
in $H^{1}(\mathbb{D})$. Also, we compute
$\|{F_{0}F+f_{0}-F_{0}K_{n}-k_{n}}\|_{1}$ $\leq\|{F_{0}F-F_{0}K_{n}}\|_{1}+\|{f_{0}-k_{n}}\|_{1}$
$\leq\|{F_{0}}\|_{2}\|{F-K_{n}}\|_{2}+\|{f_{0}-k_{n}}\|_{2}$
$\longrightarrow 0.$
This along with $\|{f-F_{0}K_{n}-k_{n}}\|_{1}\rightarrow 0$ implies that $f=F_{0}F+f_{0}$ in $H^{1}(\mathbb{D})$. But $f\in H^{2}$, and then, finally we get the desired representation of $f\in\mathcal{M}$ as
$f=F_{0}F+f_{0},$ (2.7)
along with the norm condition
$\|{f}\|^{2}=\|F\|^{2}+\|f_{0}\|^{2}.$ (2.8)
We now prove that the representation of $f$ in (2.7) under the condition (2.8) is unique. Suppose $f=F_{0}\tilde{F}+\tilde{f}_{0}=F_{0}\hat{F}+\hat{f}_{0}$ for some $\tilde{F},\hat{F}\in H^{2}_{\mathbb{C}^{p}}$ and $\tilde{f}_{0},\hat{f}_{0}\in H^{2}$. Then
$F_{0}(\tilde{F}-\hat{F})+(\tilde{f}_{0}-\hat{f}_{0})=0\in\mathcal{M},$
and by (2.8), it follows that
$\|\tilde{F}-\hat{F}\|^{2}+\|\tilde{f}_{0}-\hat{f}_{0}\|^{2}=0.$
Therefore, $\tilde{F}=\hat{F}$ and $\tilde{f}_{0}=\hat{f}_{0}$, and hence, under the norm condition (2.8), the representations of elements of $\mathcal{M}$ as in (2.7) are unique. In view of this, now we define the closed subspace $\mathcal{K}$ of $H^{2}_{\mathbb{C}^{p}}\oplus H^{2}\cong H^{2}_{\mathbb{C}^{p+1}}$ as follows:
$\mathcal{K}:=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|{f_{0}}\|^{2}=\|(F,f_{0})\|^{2}\}.$
We claim that $\mathcal{K}$ is invariant under $(T_{\varphi}\otimes I_{\mathbb{C}^{p+1}})^{*}$, that is, $(T_{\varphi}^{*}\otimes I_{\mathbb{C}^{p+1}})\mathcal{K}\subseteq\mathcal{K}$. To see this, pick $(F,f_{1})\in\mathcal{K}$. By the construction of $\mathcal{K}$, there exists $g\in\mathcal{M}$ such that
$g=F_{0}F+f_{0}.$
By the uniqueness part of $g$ as in (2.7) under the condition (2.8), it follows that $F(z)=\sum_{i=0}^{\infty}A_{i}\varphi^{i}$, and $f_{1}(z)=\sum_{j=0}^{\infty}(P_{\mathcal{K}_{\varphi}}g_{j})\varphi^{j}$. In the following, we shall identify $H^{2}$ functions with their radial limit representations. From this point of view, it follows that $\varphi\bar{\varphi}=1$, and hence
$f$ $=F_{0}F+f_{0}$
$=F_{0}F(0)+F_{0}(F-F(0))+(f_{0}-P_{\mathcal{K}_{\varphi}}f_{0})+P_{\mathcal{K}_{\varphi}}f_{0}$
$=F_{0}K_{0}(0)+F_{0}\varphi\bar{\varphi}\left(F-F(0)\right)+\varphi\bar{\varphi}\left(f_{0}-P_{\mathcal{K}_{\varphi}}f_{0}\right)+P_{\mathcal{K}_{\varphi}}f_{0}$
$=F_{0}F(0)+\varphi(F_{0}\bar{\varphi}\left(F-F(0)\right)+\bar{\varphi}\left(f_{0}-P_{\mathcal{K}_{\varphi}}f_{0}\right))+P_{\mathcal{K}_{\varphi}}f_{0}.$
From the representations of $F$ and $f_{0}$, it is easy to see that $T_{\Phi}^{*}F=\bar{\varphi}(F-F_{0}(0))$ and $T_{\varphi}^{*}f_{0}=\bar{\varphi}\left(f_{0}-P_{\mathcal{K}_{\varphi}}f_{0}\right)$, and hence
$f=FF(0)+\varphi(FT_{\Phi}^{*}F+T_{\varphi}^{*}f_{0})+P_{\mathcal{K}_{\varphi}}f_{0}.$
But $F(0)=A_{0}$, and then
$f=FA_{0}+\varphi(F_{0}T_{\Phi}^{*}F+T_{\varphi}^{*}f_{0})+P_{\mathcal{K}_{\varphi}}f_{0}.$
Recall from the above construction that $f=F_{0}A_{0}\oplus(l_{1}\varphi+P_{\mathcal{K}_{\varphi}}g_{1})$. Since
$P_{\mathcal{K}_{\varphi}}f_{0}=P_{\mathcal{K}_{\varphi}}(\sum_{j=1}^{\infty}(P_{\mathcal{K}_{\varphi}}g_{j})\varphi^{j-1})=P_{\mathcal{K}_{\varphi}}g_{1},$
we conclude that
$f=F_{0}A_{0}\oplus(l_{1}\varphi+P_{\mathcal{K}_{\varphi}}f_{0}).$
Comparing this with the above representation of $f$, we obtain (recall that $\varphi\bar{\varphi}=1$) $F_{0}T_{\Phi}^{*}F+T_{\varphi}^{*}f_{0}=l_{1}\in\mathcal{M}$. In other words
$((T_{\varphi}^{*}\otimes I_{\mathbb{C}^{p}})F,T_{\varphi}^{*}f_{0})=(T_{\Phi}^{*}F,T_{\varphi}^{*}f_{0})\in\mathcal{K},$
and hence, $\mathcal{K}$ is invariant under $T_{\varphi}^{*}\otimes I_{\mathbb{C}^{p+1}}$. This completes the proof of the theorem. ∎
Now we apply Theorem 2.1 to the case $\theta(z)=z$ for all $z\in\mathbb{D}$. We already pointed out that the $(T_{z}^{*}\otimes I_{\mathbb{C}^{p}})$-invariant subspaces are precisely model spaces (see (2.4)). Therefore, we have the following:
Corollary 2.2.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ such that
$\mathcal{M}=[F_{0},1]_{1\times(p+1)}\mathcal{K}_{\Psi},$
where $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|f_{0}\|^{2}=\|F_{0}F+f_{0}\|^{2}\}.$
In the following section, we will also talk about the opposite direction of the above result.
3. Perturbations of $T_{z}^{*}$
The setting of Theorem 2.1 is exceedingly broad, embracing invariant subspaces of finite rank perturbations of inner function-based Toeplitz operators. This one-sided result fits Hitt and Sarason’s perspectives. Indeed, their results were only one-directional. In this section, we depart from this pattern and try to explain a converse direction. Our converse, however, makes better sense for the particular case of the special inner function: $\varphi(z)=z$, $z\in\mathbb{D}$. We keep on assuming that $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets of vectors in $H^{2}$.
To address the reverse direction, we first note a consequence of Corollary 2.2: There exists a unitary operator $\mathcal{J}:\mathcal{K}\to\mathcal{M}$, such that
$\mathcal{J}(k_{1},\ldots,k_{p+1})=f_{1}k_{1}+\cdots+f_{p}k_{p}+k_{p+1},$
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$. This readily follows from the fact that an element $(F,f_{1})\in\mathcal{K}_{\Psi}$ can be written as $(F,f_{0})=(k_{1},\ldots,k_{p},k_{p+1})$ (here we are rewriting $f_{0}=k_{p+1}$ for simplicity), so that
$F_{0}F+f_{0}=f_{1}k_{1}+\cdots+f_{p}k_{p}+k_{p+1}.$
We also have the identity that $\|F\|^{2}+\|f_{0}\|^{2}=\|(F,f_{0})\|^{2}$.
We fix a rank-$m$ perturbation of $T_{z}^{*}$ as $T_{z}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, and set
$\mathcal{W}:=span\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}.$
Given a natural number $p$ (the number $p$ will be clear from the context), recall the vector-valued Hardy space identification
$H^{2}_{\mathbb{C}^{p}}\oplus H^{2}\cong H^{2}_{\mathbb{C}^{p+1}}.$
We now move on to representing constant functions in $H^{2}_{\mathbb{C}^{p+1}}$ (like the constant function $1$ in $H^{2}$). Clearly, a total of $(p+1)$ basic functions of this type exist. We represent them as $\{e_{i}\}_{i=1}^{p+1}$. In other words, for every $i=1,\ldots,p+1$, we have
$e_{i}=(0,\ldots,0,1,0,\ldots,0),$
with $1$ at the $i$-th position and zero elsewhere. We follow the outcome of Corollary 2.2. For $F=e_{i}$ with $i=1,\ldots,p$, it follows that
$F_{0}e_{i}+0=f_{i}\in\mathcal{M},$
as $\{f_{j}\}_{j=1}^{p}\subseteq\mathcal{M}$. This and
$\|e_{i}\|^{2}+\|0\|^{2}=\|(e_{i},0)\|^{2},$
ensure that
$\{e_{j}\}_{j=1}^{p}\subseteq\mathcal{K}_{\Psi}.$ (3.1)
In this context, the first part of the result below is a particular application of Corollary 2.2.
Theorem 3.1.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}u_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ such that
$\mathcal{M}=[F_{0},1]_{1\times(p+1)}\mathcal{K}_{\Psi},$ (3.2)
where $F_{0}=[f_{1},\ldots,f_{p}]$ and $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\mathcal{W}$. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|{F}\|^{2}+\|{f_{0}}\|^{2}=\|F_{0}F+f_{0}\|^{2}\},$ (3.3)
and there exists a unitary operator $\mathcal{J}:\mathcal{K}_{\Psi}\to\mathcal{M}$ such that
$\mathcal{J}(k_{1},\ldots,k_{p+1})=f_{1}k_{1}+\cdots+f_{p}k_{p}+k_{p+1},$ (3.4)
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$. Conversely, if $\mathcal{M}$ has the representation (3.2) for some orthonormal basis $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ with $\mathcal{K}_{\Psi}$ as in (3.3), and a unitary $\mathcal{J}$ as in (3.4), then $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i}$.
Proof.
We only need to prove the sufficient part. In this case, note that
=$\mathcal{M}=\{f\in H^{2}:f=F_{0}F+f_{0}\text{ with }\|f\|^{2}=\|F\|^{2}+\|{f_{0}}\|^{2}\text{ for some }(F,f_{0})\in\mathcal{K}_{\Psi}\}.$
Given $f\in\mathcal{M}$, the above representation of $\mathcal{M}$ implies the existence of $(F,f_{0})\in\mathcal{K}_{\Psi}$ such that $f=F_{0}F+f_{0}$. If we write $F=(k_{0}^{1},\ldots,k_{0}^{p})^{t}\in H^{2}_{\mathbb{C}^{p}}$, then
$(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i})f$ $=[T_{z}^{*}-\sum_{i=1}^{p}\langle\cdot,f_{i}\rangle T_{z}^{*}f_{i}](F_{0}F+f_{0})$
$=\frac{F_{0}(z)F(z)-F_{0}(0)F(0)}{z}+T_{z}^{*}f_{0}-\sum_{i=1}^{p}\langle F_{0}F+f_{0},f_{i}\rangle T_{z}^{*}f_{i}.$
On the other hand
$\mathcal{J}(k_{0}^{1},\ldots,k_{0}^{p},1)=F_{0}F+f_{0},$
and
$\mathcal{J}e_{i}=f_{i},$
for all $i=1,\ldots,p$, where $e_{i}=(0,\ldots,0,1,0,\ldots,0)$ with $1$ at the $i$-th position and zero elsewhere. This conclusion arises immediately from the identity (3.1). Since $\mathcal{J}$ is an isometry (in fact, unitary), for each $i=1,\ldots,p$, it follows that
$\begin{split}\langle F_{0}K_{0}+f_{0},f_{i}\rangle&=\langle\mathcal{J}(k_{0}^{1},\ldots,k_{0}^{p},1),\mathcal{J}e_{i}\rangle\\ &=\langle(k_{0}^{1},\ldots,k_{0}^{p},1),e_{i}\rangle\\ &=k_{0}^{i}(0).\end{split}$
Moreover, since $T_{z}^{*}F=\frac{F-F(0)}{z}$ and $T_{z}^{*}F_{0}=\frac{F_{0}-F_{0}(0)}{z}$, we have
$(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i})f$ $=F_{0}T_{z}^{*}(F)+T_{z}^{*}(F_{0})F(0)+T_{z}^{*}f_{0}-\sum_{i=1}^{p}k_{0}^{i}(0)T_{z}^{*}f_{i}$
$=F_{0}(T_{z}^{*}F)+(T_{z}^{*}F_{0})F(0)+T_{z}^{*}f_{0}-(T_{z}^{*}F_{0})F(0)$
$=F_{0}(T_{z}^{*}F)+T_{z}^{*}f_{0}$
$\in\mathcal{M},$
proving that $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i}$. ∎
Considering the converse half of the previous statement, it is evident that the theorem concerns the classification of closed subspaces of $H^{2}$ that are invariant under Sarason-type perturbations of the backward shift $T_{z}^{*}$. Furthermore, in the setting of Corollary 2.2, the same proof of the converse part also shows that $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i}$. The following alternative can be considered a variation of Corollary 2.2:
Corollary 3.2.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ such that
$\mathcal{M}=[F_{0},1]_{1\times(p+1)}\mathcal{K}_{\Psi},$
where $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}u_{1},\ldots,P_{\mathcal{M}}u_{m}\}$. Moreover, $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i}$.
4. Perturbations of $T_{z}$
Let $T\in\mathcal{B}(\mathcal{H})$ and let $\mathcal{M}$ is a closed subspace of $\mathcal{H}$. In the introductory section, we already pointed out the elementary fact that $T\mathcal{M}\subseteq\mathcal{M}$ if and only if $T^{*}\mathcal{M}^{\perp}\subseteq\mathcal{M}^{\perp}$. However, if $\mathcal{M}$ is explicit, then it need not be the case for $\mathcal{M}^{\perp}$. For instance, the concrete representations of finite-rank perturbations of $T_{z}^{*}$-invariant subspaces do not seem to yield concrete representations of finite-rank perturbations of $T_{z}$-invariant subspaces. In this section, we examine a class of perturbations of $T_{z}$, and obtain explicit representations of invariant subspaces of the perturbations. As hinted above, the primary method will involve exploiting the adjoint of the operator used in Theorem 2.1.
As usual, we consider orthonormal sets of vectors $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ in $H^{2}$. Given a closed subspace $\mathcal{M}\subseteq H^{2}$, we let
$\mathcal{L}:=span\{P_{\mathcal{M}^{\perp}}v_{1},\ldots,P_{\mathcal{M}^{\perp}}v_{m}\},$
and assume that
$p=\dim\mathcal{L}.$
In comparison to Theorem 2.1, the following assumes that there is an orthonormal basis for $\mathcal{L}$ that is contained in $H^{\infty}$.
Theorem 4.1.
Let $\mathcal{M}$ be an invariant subspace of $T_{z}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$. Suppose $\{\varphi_{i}\}_{i=1}^{p}\subseteq H^{\infty}$ is an orthonormal basis for $\mathcal{L}$. Then
=$\mathcal{M}=\{g\in H^{2}:(T_{\bar{\varphi}_{1}}g,\ldots,T_{\bar{\varphi}_{p}}g,g)\in\Psi H^{2}_{\mathbb{C}^{p^{\prime}}}\},$
for some $p^{\prime}\leq p+1$ and inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$.
Proof.
Note that $\mathcal{M}^{\perp}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}u_{i}\otimes v_{i}$. By Theorem 2.1, there exist $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
$\mathcal{M}^{\perp}=[\varphi_{1},\ldots,\varphi_{p},1]\mathcal{K}_{\Psi}.$
Let us assume that $g\in H^{2}$. Then $g\in\mathcal{M}^{\perp}$ if and only if $\langle g,f\rangle=0$ for all $f\in\mathcal{M}^{\perp}$. By the above representation of $\mathcal{M}^{\perp}$, we consider an arbitrary $f$ from $\mathcal{M}^{\perp}$ as $f=\sum_{i=1}^{p}\varphi_{i}k_{i}+k_{p+1}$ for some (again, arbitrary) $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$, and compute
$\langle g,f\rangle$ $=\langle g,\sum_{i=1}^{p}\varphi_{i}k_{i}+k_{p+1}\rangle$
$=\langle g,\sum_{i=1}^{p}\varphi_{i}k_{i}\rangle+\langle g,k_{p+1}\rangle$
$=\sum_{i=1}^{p}\langle T_{\bar{\varphi}_{i}}g,k_{i}\rangle+\langle g,k_{p+1}\rangle.$
Therefore, $\langle g,f\rangle=0$ for all $f\in\mathcal{M}^{\perp}$ if and only if
$(T_{\bar{\varphi}_{1}}g,\ldots,T_{\bar{\varphi}_{p}}g,g)\in\Psi H^{2}_{\mathbb{C}^{p^{\prime}}}(\mathbb{D}),$
completing the proof of the theorem. ∎
To formulate a more convenient converse statement, we will now concentrate on Sarason-type perturbations of the forward shift (see (1.6)). Recall, for each $\varphi\in L^{\infty}$, the Toeplitz operator $T_{\varphi}$ on $H^{2}$ is defined by
$T_{\varphi}f=P_{H^{2}}(\varphi f),$
for all $f\in H^{2}$. We will need the well-known algebraic characterization of Toeplitz operators: Let $T\in\mathcal{B}(H^{2})$. Then $T$ is Toeplitz if and only if
$T_{z}^{*}TT_{z}=T.$ (4.1)
Theorem 4.2.
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. Assume that $\{\varphi_{i}\}_{i=1}^{p}\subseteq H^{\infty}$ is an orthonormal basis for $\mathcal{L}$. If $\mathcal{M}$ is invariant under $T_{z}-\sum_{i=1}^{m}T_{z}^{*}u_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
=$\mathcal{M}=\{g\in H^{2}:(T_{\bar{\varphi}_{1}}g,\ldots,T_{\bar{\varphi}_{p}}g,g)\in\Psi H^{2}_{\mathbb{C}^{p^{\prime}}}\},$
and the operator $\mathcal{J}:\mathcal{K}_{\Psi}\to\mathcal{M}^{\perp}$ defined by
$\mathcal{J}(k_{1},\ldots,k_{p+1})=\varphi_{1}k_{1}+\cdots+\varphi_{p}k_{p}+k_{p+1},$
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$ is unitary. Conversely, if $\mathcal{M}$ admits the above representation along with the existence of the unitary $\mathcal{J}$, then $\mathcal{M}$ is invariant under $T_{z}-\sum_{i=1}^{p}\varphi_{i}\otimes T_{z}^{*}\varphi_{i}$.
Proof.
We only have to prove the reverse direction. Note that $T_{\varphi_{i}}^{*}=T_{\bar{\varphi}_{i}}$ for all $i=1,\ldots,p$, and hence
$\mathcal{J}^{*}h=(T_{\bar{\varphi}_{1}}h,\ldots,T_{\bar{\varphi}_{p}}h,h)\in\mathcal{K}_{\Psi},$
for all $h\in\mathcal{M}^{\perp}$. Let $g\in\mathcal{M}$. Then
$(T_{z}-\sum_{i=1}^{p}\varphi_{i}\otimes T_{z}^{*}\varphi_{i})g=T_{z}g-\tilde{g}$
where $\tilde{g}=\sum_{i=1}^{p}\langle g,T_{z}^{*}\varphi_{i}\rangle\varphi_{i}$, and also $(T_{\bar{\varphi}_{1}}g,\ldots,T_{\bar{\varphi}_{p}}g,g)\in\Psi H^{2}_{\mathbb{C}^{p^{\prime}}}(\mathbb{D})$. Hence
$(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)(T_{z}g-\tilde{g})$ $=(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)T_{z}g-(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)\tilde{g}$
$=(T_{\bar{\varphi}_{1}}T_{z},\ldots,T_{\bar{\varphi}_{p}}T_{z},T_{z})g-(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)\tilde{g}$
$=(T_{z}T_{z}^{*}T_{\bar{\varphi}_{1}}T_{z},\ldots,T_{z}T_{z}^{*}T_{\bar{\varphi}_{p}}T_{z},T_{z})g-(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)\tilde{g}$
$\quad+(P_{\mathbb{C}}T_{\bar{\varphi}_{1}}T_{z},\ldots,P_{\mathbb{C}}T_{\bar{\varphi}_{p}}T_{z},0)g,$
as $T_{z}T_{z}^{*}+P_{\mathbb{C}}=I$, where $P_{\mathbb{C}}$ denotes the orthogonal projection onto the one-dimensional space of constant functions in $H^{2}$. Since $T_{\bar{\varphi}_{j}}$ is a Toeplitz operator, $T_{z}^{*}T_{\bar{\varphi}_{p}}T_{z}=T_{\bar{\varphi}_{p}}$ for all $j=1,\ldots,p$ (see (4.1)), and consequently
$(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)(T_{z}g-\tilde{g})$ $=T_{z}(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)g-\mathcal{J}^{*}(\sum_{i=1}^{p}\langle g,T_{z}^{*}\varphi_{i}\rangle\mathcal{J}e_{i})+\sum_{j=1}^{p}\langle T_{z}g,\varphi_{j}\rangle e_{j}$
$=T_{z}(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)g-\sum_{i=1}^{p}\langle g,T_{z}^{*}\varphi_{i}\rangle e_{i}+\sum_{j=1}^{p}\langle g,T_{z}^{*}\varphi_{j}\rangle e_{j}$
$=T_{z}(T_{\bar{\varphi}_{1}},\ldots,T_{\bar{\varphi}_{p}},I)g$
$\in\Psi H^{2}_{\mathbb{C}^{p^{\prime}}}(\mathbb{D}),$
which implies that $\mathcal{M}$ is invariant under the Sarason-type perturbation $T_{z}-\sum_{i=1}^{p}\varphi_{i}\otimes T_{z}^{*}\varphi_{i}$. ∎
Now let us reexamine the discussion that takes place at the end of Section 3. In particular, we continue the results described in Corollary 3.2. Along this line, we note that the subspace $\mathcal{M}$ above is also invariant under $T_{z}-\sum_{i=1}^{m}\varphi_{i}\otimes T_{z}^{*}\varphi_{i}$. To confirm this, we proceed as follows: Pick $f\in\mathcal{M}^{\perp}$. By Theorem 2.1, we have
$f=F_{0}F+f_{0},$
where $F=\sum_{i=0}^{\infty}A_{i}z^{i}$, $A_{i}\in\mathbb{C}^{p}$ for all $i$, and $f_{0}=\sum_{i=0}^{\infty}g_{i}(0)z^{i}$. Then
$\begin{split}(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})f&=(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})(F_{0}F+f_{0})\\ &=T_{z}^{*}(F_{0}F)+T_{z}^{*}f_{0}-\sum_{i=1}^{m}\langle F_{0}F+f_{0},\varphi_{i}\rangle T_{z}^{*}\varphi_{i}.\end{split}$
Now assume $F=(k_{0}^{1},\ldots,k_{0}^{p})$, and proceed with the proof of Theorem 3.1. We find
$\langle F_{0}K_{0}+f_{0},f_{i}\rangle=k_{0}^{i}(0),$
and finally
$(T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}\varphi_{i}\otimes\varphi_{i})f=F_{0}(T_{z}^{*}F)+T_{z}^{*}f_{0}\in\mathcal{M}^{\perp},$
which proves that $\mathcal{M}^{\perp}$ is invariant under the Sarason-type perturbation $T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}\varphi_{i}\otimes\varphi_{i}$.
5. Almost invariant subpsaces
First, we recall the definition of almost invariant subspaces (see [4]). Let $T\in\mathcal{B}(\mathcal{H})$, and let $\mathcal{M}\subseteq\mathcal{H}$ be a closed subspace. Then $\mathcal{M}$ is almost invariant for $T$ if there exists a finite-dimensional subspace $\mathcal{F}\subseteq\mathcal{H}$ such that
$T\mathcal{M}\subseteq\mathcal{M}\oplus\mathcal{F}.$
Recall also that the defect of $\mathcal{M}$ is the lowest possible dimension of such a space $\mathcal{F}$. This section aims to link almost invariant subspaces of $T_{z}^{*}$ with the invariant subspaces of finite-rank perturbations for $T_{z}^{*}$. The link is quite natural and could even be considered a simple observation. However, this link in particular gives a new insight into the theory of almost invariant subspaces and unifies the perturbation theory from an invariant subspace stance. We present the result in finer generality. We will assume $\mathcal{H}$ is a Hilbert space, and $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets in $\mathcal{H}$.
Theorem 5.1.
Let $T\in\mathcal{B}(\mathcal{H})$ and let $\mathcal{M}\subseteq\mathcal{H}$ be a closed subspace. If $\mathcal{M}\subseteq\mathcal{H}$ is invariant under $T-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then $\mathcal{M}$ is almost invariant under $T$ with defect at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T$ with defect $m$, then $\mathcal{M}$ is invariant under $T-\sum_{i=1}^{m}f_{i}\otimes T^{*}f_{i}$, where $\{f_{i}\}_{i=1}^{m}$ is an orthonormal basis for the defect space.
Proof.
Suppose $\mathcal{M}\subseteq\mathcal{H}$ is invariant under $T-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, that is
$(T-\sum_{i=1}^{m}v_{i}\otimes u_{i})\mathcal{M}\subseteq\mathcal{M}.$
Note that
$(v_{i}\otimes u_{i})\mathcal{M}=\{\langle f,u_{i}\rangle v_{i}:f\in\mathcal{M}\}=\mathbb{C}v_{i},$
for all $i=1,\ldots,m$. Therefore
$\begin{split}T\mathcal{M}&=\Big{(}(T-\sum_{i=1}^{m}v_{i}\otimes u_{i})+\sum_{i=1}^{m}v_{i}\otimes u_{i}\Big{)}\mathcal{M}\\ &\subseteq\mathcal{M}+\text{span}\{v_{i}:i=1,\ldots,m\},\end{split}$
and hence, $\mathcal{M}$ is almost invariant under $T$ with defect space $\subseteq\text{span}\{v_{i}:i=1,\ldots,m\}$. This also says, in particular, that the defect dimension is at most $m$. For the converse direction, assume that $\mathcal{M}$ is almost invariant under $T$ with defect $m$ and that $\mathcal{F}$ is the $m$-dimensional defect space having an orthonormal basis, say $\{f_{i}:1\leq i\leq m\}$. Therefore, we have $T\mathcal{M}\subseteq\mathcal{M}\oplus\mathcal{F}$. Fix $f\in\mathcal{M}$. Since $Tf\in\mathcal{M}\oplus\mathcal{F}$, there exist $f_{\mathcal{M}}\in\mathcal{M}$ and $f_{\mathcal{F}}\in\mathcal{F}$ such that
$Tf=f_{\mathcal{M}}\oplus f_{\mathcal{F}}.$
Also note that $f_{\mathcal{F}}=\sum_{i=1}^{m}\langle f_{\mathcal{F}},f_{i}\rangle f_{i}$. Therefore, we have
$\begin{split}(T-\sum_{i=1}^{m}f_{i}\otimes T^{*}f_{i})f&=f-\sum_{i=1}^{m}\langle f,T^{*}f_{i}\rangle f_{i}\\ &=f-\sum_{i=1}^{m}\langle f_{\mathcal{M}}\oplus f_{\mathcal{F}},f_{i}\rangle f_{i}\\ &=f-\sum_{i=1}^{m}\langle f_{\mathcal{F}},f_{i}\rangle f_{i}\\ &=(f_{\mathcal{M}}\oplus f_{\mathcal{F}})-f_{\mathcal{F}}\\ &=f_{\mathcal{M}}\\ &\in\mathcal{M}.\end{split}$
This proves that $\mathcal{M}$ is invariant under $T-\sum_{i=1}^{m}f_{i}\otimes T^{*}f_{i}$. ∎
In particular, we have the following in the setting of $T^{*}_{z}$: Let $\mathcal{M}$ be a closed subspace of $H^{2}$, and let $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ be orthonormal sets in $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect dimension at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect dimension $m$, then $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{m}f_{i}\otimes T_{z}f_{i}$, where $\{f_{i}:1\leq i\leq m\}$ is an orthonormal basis of the $m$ dimensional defect space.
In Theorem 3.1, we have already classified the invariant subspaces of finite rank perturbations of $T_{z}^{*}$ of the above type. As a result, the above observation, along with Theorem 3.1, easily implies the following classification of almost invariant subspaces of $T_{z}^{*}$:
Corollary 5.2.
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. If $\mathcal{M}$ is almost invariant under $T_{z}^{*}$, then there exist natural numbers $p^{\prime}<p$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$, and a unitary operator $\mathcal{J}:\mathcal{K}_{\Psi}\to\mathcal{M}$ such that
$\mathcal{M}=[F_{0},1]_{1\times(p+1)}\mathcal{K}_{\Psi},$ (5.1)
where $\{f_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}(u_{1}),\ldots,P_{\mathcal{M}}(u_{m})\}$, and $\{u_{1},\ldots,u_{m}\}$ is an orthonormal basis for the defect space, and
$\mathcal{J}(k_{1},\ldots,k_{p+1})=f_{1}k_{1}+\cdots+f_{p}k_{p}+k_{p+1},$ (5.2)
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|{F}\|^{2}+\|{f_{0}}\|^{2}=\|F_{0}F+f_{0}\|^{2}\}.$ (5.3)
Conversely, if $\mathcal{M}$ has the representation (5.1) for some orthonormal set of vectors $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ with $\mathcal{K}_{\Psi}$ as in (5.3), and a unitary $\mathcal{J}$ as in (5.2), then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$.
A similar classification of almost invariant subspaces of $T_{z}^{*}$ was earlier obtained by Chalendar, Gallardo-Gutiérrez, and Partington [4, Theorem 3.4]. However, the present result differs substantially. First, it connects with a topic that appears to have a different flavor, namely finite-rank perturbations of $T_{z}^{*}$; second, this comes as a byproduct of finite-rank perturbations of $T_{z}^{*}$. Third, the connection with the perturbation theory made the representations of almost invariant subspaces more explicit. For instance, the representation of the model space $\mathcal{K}_{\Psi}$ is a new addition compared to the one derived in [4, Theorem 3.4].
6. Nearly invariant subspaces
The aim of this section is to establish a connection between our previous results on finite rank perturbations and the concept of nearly invariant subspaces. At first, we set up the notations. Let $\{z_{k}\}_{k=1}^{n}$ be $n$-points (repetition is allowed) in $\mathbb{D}$. Assume that $z_{j}=0$ for some $j=1,\ldots,n$. Define the corresponding Blaschke product $B_{n}$ by
$B_{n}(z)=\prod_{k=1}^{n}\dfrac{z_{k}-z}{1-\bar{z}_{k}z},$
for all $z\in\mathbb{D}$. Clearly, $B_{n}\in H^{\infty}$ is an inner function, and $B_{n}(0)=0$. Then the corresponding model space $\mathcal{K}_{B_{n}}=H^{2}/B_{n}H^{2}$ is $n$-dimensional. Throughout this section, we will fix the Blaschke product mentioned above. We recall the following from Definition 1.4: A closed subspace $\mathcal{M}\subseteq H^{2}$ is nearly $T^{*}_{z,B_{n}}$-invariant if $T_{z}^{*}f\in\mathcal{M}$ for all
$f\in\mathcal{M}\cap B_{n}H^{2}.$
If $n=1$, then $B_{1}(z)=z$, and hence, nearly $T^{*}_{z,B_{1}}$-invariant subspaces are precisely nearly $T_{z}^{*}$-invariant. For the basic preparation, we recall that $H^{2}$ is a reproducing kernel Hilbert space corresponding to the Szegö kernel
$k(z,w)=\frac{1}{1-\bar{w}z},$
for all $z,w\in\mathbb{D}$. The kernel function $k_{w}$ at $w\in\mathbb{D}$ is defined by
$k_{w}(z)=k(z,w),$
for all $z\in\mathbb{D}$. We know that the set of kernel functions $\{k_{w}:w\in\mathbb{D}\}$ is total in $H^{2}$ and also satisfy the reproducing property
$f(w)=\langle f,k_{w}\rangle,$
for all $f\in H^{2}$ and $w\in\mathbb{D}$. We need the following lemma, which involves Sarason-type perturbations:
Lemma 6.1.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}}$, then
$T_{z}^{*}(\mathcal{M}\cap B_{n}H^{2})\subseteq\mathcal{M}.$
Proof.
Let $f\in\mathcal{M}\cap B_{n}H^{2}$. There exists $g\in H^{2}$ such that $f=B_{n}g$. Then
$(T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}})f=\langle f,k_{z_{i}}\rangle T_{z}^{*}k_{z_{i}}=\langle B_{n}g,k_{z_{i}}\rangle T_{z}^{*}k_{z_{i}}=B_{n}(z_{i})g(z_{i})T_{z}^{*}k_{z_{i}}=0,$
as $B_{n}(z_{i})=0$ for all $i=1,\ldots,n$. By assumption, we have $(T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}})f\in\mathcal{M}$, and hence
$T_{z}^{*}f=(T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}})f+(\sum_{i=1}^{n}T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}})f=(T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}k_{z_{i}}\otimes k_{z_{i}})f\in\mathcal{M},$
completes the proof of the lemma. ∎
We need another lemma:
Lemma 6.2.
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. Then
$\dim\left(\mathcal{M}\ominus(\mathcal{M}\cap B_{n}H^{2})\right)\leq m.$
Proof.
Recall that the dimension of the model space $\mathcal{K}_{B_{n}}=H^{2}\ominus B_{n}H^{2}$ is $n$. Suppose $\{f_{1},\ldots,f_{n}\}$ is an orthonormal basis for $\mathcal{K}_{B}$. Set
$\mathcal{L}=span\{P_{\mathcal{M}}f_{1},\ldots,P_{\mathcal{M}}f_{n})\}.$
Clearly, $\mathcal{L}\subseteq\mathcal{M}$. Now, for each $f\in\mathcal{M}\cap B_{n}H^{2}$ and $j=1,\ldots,n$, we have
$\langle P_{\mathcal{M}}f_{j},f\rangle=\langle f_{j},f\rangle=0,$
as $f_{j}\perp B_{n}H^{2}$. This implies that $\mathcal{L}\subseteq\mathcal{M}\ominus(\mathcal{M}\cap B_{n}H^{2})$. We claim that this inclusion is part of equality. To see this, pick $f\in\mathcal{M}$ and assume that $\langle f,P_{\mathcal{M}}f_{j}\rangle=0$ for all $j=1,\ldots,n$. This is equivalent to say that $\langle f,f_{j}\rangle=0$ for all $j=1,\ldots,n$, and hence $f\in B_{n}H^{2}$. Thus $f\in\mathcal{M}\cap B_{n}H^{2}$, and so $\mathcal{L}=\mathcal{M}\ominus(\mathcal{M}\cap B_{n}H^{2})$, completing the proof of the lemma. ∎
We are now ready to prove the invariant subspace theorem. It is of importance to point out that the proof follows the same structure as Theorem 2.1. Hence, the subsequent proof will be concise and will exclude details that resemble the proof of Theorem 2.1. Additionally, it is worth recalling that the proof of Theorem 2.1 is partially influenced by the methodology employed by Hitt and Sarason.
Theorem 6.3.
Let $\mathcal{M}\subseteq{H}^{2}$ be a nontrivial closed subspace of $H^{2}$. Then $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant if and only if there exist natural numbers $r^{\prime}\leq r\leq n$ and inner function $\Theta\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{r^{\prime}},\mathbb{C}^{r})}$ such that $\Theta(0)=0$ and
$\mathcal{M}=[g_{1},\ldots,g_{r}]_{1\times r}\mathcal{K}_{\Theta},$
where $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B{H}^{2})$, and
$\mathcal{J}(k_{1},\ldots,k_{r})=g_{1}k_{1}+\cdots+g_{r}k_{r},$
for all $(k_{1},\ldots,k_{r})\in\mathcal{K}_{\Theta}$, defines a unitary operator $\mathcal{J}:\mathcal{K}_{\Theta}\rightarrow\mathcal{M}$.
Proof.
Suppose $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant. We claim that
$\mathcal{M}\nsubseteq B_{n}{H}^{2}.$
Indeed, if $\mathcal{M}\subseteq B{H}^{2}$, then $\mathcal{M}\subseteq z{H}^{2}$ as $B_{n}(0)=0$. So, for any
$f=\sum_{n=0}^{\infty}a_{n}z^{n}\in\mathcal{M},$
we have $a_{0}=0$. By the definition of nearly $T^{*}_{z,B_{n}}$-invariant, we have $T_{z}^{*}f\in\mathcal{M}$, which gives $\sum_{n=1}^{\infty}a_{n}z^{n-1}\in\mathcal{M}\subseteq z{H}^{2}$. But now $a_{1}=0$, and hence, in a similar way, we conclude that $a_{n}=0$ for all $n\geq 0$. Therefore, $f=0$, and hence, =$\mathcal{M}=\{0\}$. This proves the claim, and consequently, we have
$\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})\neq\{0\}.$
This and Lemma 6.2 then implies that
$r:=\dim(\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2}))\in[1,m].$
Suppose $\{g_{1},\ldots,g_{r}\}$ be an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})$. We decompose $\mathcal{M}$ as
$\mathcal{M}=\big{(}\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})\big{)}\oplus\big{(}\mathcal{M}\cap B_{n}{H}^{2}\big{)}.$
Fix $f\in\mathcal{M}$, and write
$f=f_{0}\oplus h_{1}$
for some $f_{0}\in\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})$ and $h_{1}\in\mathcal{M}\cap B_{n}{H}^{2}$. Assume that $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})$, and set $G_{0}=[g_{1},\ldots,g_{r}]_{1\times r}$. Then
$f_{0}=a_{01}g_{1}+\cdots+a_{0r}g_{r}=G_{0}A_{0},$
where $A_{0}=[a_{01},\ldots,a_{0r}]^{t}\in\mathbb{C}^{r}$. We also have that
$\|{f}\|^{2}=\|{A_{0}}\|^{2}+\|{h_{1}}\|^{2}.$
Since $h_{1}\in\mathcal{M}\cap B_{n}{H}^{2}$, by the definition of nearly $T^{*}_{z,B_{n}}$-invariant subspace, $T_{z}^{*}h_{1}\in\mathcal{M}$, where, on the other had, $B_{n}(0)=0$ implies $h_{1}=zf_{1}$ for some $f_{1}\in\mathcal{M}$. Then
$f=G_{0}A_{0}\oplus zf_{1},$
and $\|{f}\|^{2}=\|{A_{0}}\|^{2}+\|{f_{1}}\|^{2}$. We are now in exactly the same situation as in the proof of Theorem 2.1, specifically the identity obtained in (2.5). Thus, we take the argument further and conclude similarly that
$f=G_{0}K,$
with $\|{f}\|^{2}=\|{K}\|^{2}$, where, $K=\sum_{n=0}^{\infty}A_{n}z^{n}\in H^{2}_{\mathbb{C}^{r}}$. With this in mind, define
$\mathcal{K}:=\left\{K\in H^{2}_{\mathbb{C}^{r}}:G_{0}K\in\mathcal{M},\text{ and }\|G_{0}K\|=\|{K}\|\right\}\subseteq H^{2}_{\mathbb{C}^{r}}.$
Pick $K\in\mathcal{K}$. Then there exists $f\in\mathcal{M}$ such that $f=G_{0}K$. By the above construction, we know $f=G_{0}A_{0}\oplus zf_{1}$ with $A_{0}=K(0)$. Then
$G_{0}\left(\frac{K(z)-K(0)}{z}\right)=f_{1}\in\mathcal{M},$
implying $T_{z}^{*}K\in\mathcal{K}$. Therefore by the Beurling-Lax theorem, there exist a natural number $r^{\prime}\leq r$ and an inner function $\Theta\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{r^{\prime}},\mathbb{C}^{r})}$ such that $\mathcal{K}=\mathcal{K}_{\Theta}$. In other words, $\mathcal{M}=G_{0}\mathcal{K}_{\Theta}$. Our next aim is to show $\Theta(0)=0$. Denote by $\{e_{i}\}_{i=1}^{r}$ the standard orthonormal basis for $\mathbb{C}^{r}$. For any $i=1,\ldots,r$, if we set $K=1\otimes e_{i}$, then $G_{0}K=g_{i}$ together with $\|g_{i}\|=\|G_{0}K\|=\|e_{i}\|=1$ implies $1\otimes e_{i}\in\mathcal{K}_{\Theta}$. This means
$1\otimes\eta\in\mathcal{K}_{\Theta},$
for all $\eta\in\mathbb{C}^{r}$, and hence
$\Theta(0)^{*}\eta=T_{\Theta}^{*}(1\otimes\eta)=0,$
for all $\eta\in\mathbb{C}^{r}$. Therefore, $\Theta(0)=0$. The existence of unitary $\mathcal{J}$ follows from the description of the space $\mathcal{K}$.
For the converse direction, pick $f=G_{0}K\in\mathcal{M}$ for some $K=(k_{1},\ldots,k_{r})\in\mathcal{K}_{\Theta}$. Assume that $f\in\mathcal{M}\cap B_{n}{H}^{2}$. Since $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})$, for $i\in\{1,\ldots,r\}$, we have
$\langle f,g_{i}\rangle=\langle g_{1}k_{1}+\cdots+g_{r}k_{r},g_{i}\rangle=0,$
which implies $k_{i}(0)=0$, and hence $K(0)=0$. Now $f=G_{0}K$ implies that $f(0)=0$, and then from $T_{z}^{*}f=\frac{f-f(0)}{z}$, it follows that
$T_{z}^{*}f=\frac{G_{0}K}{z}=G_{0}\frac{K-K(0)}{z}=G_{0}(T_{z}^{*}K)\in G_{0}\mathcal{K}_{\Theta}=\mathcal{M},$
from which we conclude that $\mathcal{M}$ is $T^{*}_{z,B_{n}}$-invariant. ∎
Considering the case $B_{1}(z)=z$, we can recover the structure of classical nearly $T_{z}^{*}$-invariant subspaces, which has been described by Hitt [9] and further modified by Sarason [15]. In this case, nearly $T_{z}^{*}$-invariant subspaces are of the form $gK_{\varphi}$ where $\mathcal{K}_{\varphi}\subseteq H^{2}$ is a model space for some inner function $\varphi\in H^{\infty}$ and $g$ is an isometric multiplier on $\mathcal{K}_{\varphi}$ (for instance, see [6]).
7. Perturbed Toeplitz operators
The purpose of this section is to provide some examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. From this perspective, we remind the reader that the kernels of Toeplitz operators are nearly $T_{z}^{*}$-invariant subspaces. Along this line, we prove that the kernels of certain finite-rank perturbations of Toeplitz operators are nearly $T^{*}_{z,B_{n}}$-invariant. Consequently, in addition to the theory of finite rank perturbations of $T_{z}^{*}$, the results of this section can be considered an additional justification for the new notion of nearly $T^{*}_{z,B_{n}}$-invariant subspaces.
We adopt the setting of Section 6 and continue our discussion with nearly $T^{*}_{z,B_{n}}$-invariant subspaces. Therefore, $B_{n}$ is a fixed Blaschke product with $\{z_{k}\}_{k=1}^{n}$ as the zero set, and we assume that $B_{n}(0)=0$. Recall that
$\dim\mathcal{K}_{B_{n}}=n,$
where $\mathcal{K}_{B}$ is the model space corresponding to the inner function $B_{n}$. For $\varphi\in L^{\infty}$, the perturbed Toeplitz operator
$T^{\varphi}_{B}:=T_{\varphi}+\sum_{i=1}^{m}\langle\cdot,S^{*}(h_{i})\rangle h_{i},$
is atmost $m$-rank perturbation of the Toeplitz operator $T_{\varphi}$. Given an arbitrary orthonormal basis $\{f_{i}\}_{i=1}^{n}$ of the model space $\mathcal{K}_{B_{n}}$, we will be interested in the finite-rank perturbation of $T_{\varphi}$ as
$T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i}.$
In the following, we prove that the kernel of the above operator is nearly $T^{*}_{z,B_{n}}$-invariant subspace:
Theorem 7.1.
Let $\varphi\in L^{\infty}$. Then, for each orthonormal basis $\{f_{i}\}_{i=1}^{n}$ for the model space $\mathcal{K}_{B_{n}}$, the kernel space
$\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i}),$
is nearly $T^{*}_{z,B_{n}}$-invariant.
Proof.
Pick an arbitrary
$f\in\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})\cap B_{n}H^{2}.$
Our goal is to prove that $T_{z}^{*}f\in\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})$. Since
$(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})f=0,$
it follows that
$T_{\varphi}f+\sum_{i=1}^{n}\langle T_{z}f,f_{i}\rangle f_{i}=0.$
Since $f\in B_{n}H^{2}$, we have $T_{z}f\in B_{n}H^{2}$. Also, $\{f_{i}\}_{i=1}^{n}$ is an orthonormal basis for $\mathcal{K}_{B_{n}}$. This implies
$\sum_{i=1}^{n}\langle T_{z}f,f_{i}\rangle f_{i}=0,$
and then
$T_{\varphi}f=0.$
Next, we recall that $T_{z}T_{z}^{*}=I-P_{\mathbb{C}}$, where $P_{\mathbb{C}}$ denotes the orthogonal projection of $H^{2}$ onto the one-dimensional space of all constant functions. In the present scenario, $f$ is in $B_{n}H^{2}$ and $B_{n}(0)=0$, which implies $f(0)=0$. Therefore, $P_{\mathbb{C}}f=0$ and hence
$T_{z}T_{z}^{*}f=f.$
We use this and the Toeplitz operator identity $T_{z}^{*}T_{\varphi}T_{z}=T_{\varphi}$ (see (4.1)) to compute
$(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})T_{z}^{*}f$ $=T_{\varphi}(T_{z}^{*}f)+\sum_{i=1}^{n}\langle T_{z}^{*}f,T_{z}^{*}f_{i}\rangle f_{i}$
$=T_{z}^{*}T_{\varphi}T_{z}(T_{z}^{*}f)+\sum_{i=1}^{n}\langle T_{z}T_{z}^{*}f,f_{i}\rangle f_{i}$
$=T_{z}^{*}T_{\varphi}f+\sum_{i=1}^{n}\langle f,f_{i}\rangle f_{i}.$
Since $f\in BH^{2}$ and $\{f_{i}\}_{i=1}^{n}$ is an orthonormal basis for $\mathcal{K}_{B_{n}}$, it is clear that $\sum_{i=1}^{n}\langle f,f_{i}\rangle f_{i}=0$. Then $T_{\varphi}f=0$ implies that
$T_{z}^{*}f\in\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i}),$
and proves the claim that $\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i})$ is nearly $T^{*}_{z,B_{n}}$-invariant. ∎
Evidently, the case $B_{1}(z)=z$ yields the known fact that the kernel of a Toeplitz operator is nearly $T_{z}^{*}$-invariant. The following observation is a comparison between nearly $T_{z}^{*}$-invariant subspaces and nearly $T^{*}_{z,B_{n}}$-invariant subspaces:
Proposition 7.2.
Every nearly $T_{z}^{*}$-invariant subspace is a nearly $T^{*}_{z,B_{n}}$-invariant subspace.
Proof.
Let $\mathcal{M}$ be a nearly $T_{z}^{*}$-invariant subspace of $H^{2}$. Since $B_{n}(0)=0$, it follows that $B_{n}H^{2}\subseteq zH^{2}$, and hence
$\mathcal{M}\cap B_{n}H^{2}\subseteq\mathcal{M}\cap zH^{2}.$
Therefore, if $f\in\mathcal{M}\cap B_{n}H^{2}$, then $f\in\mathcal{M}\cap zH^{2}$, and consequently $T_{z}^{*}f\in\mathcal{M}$. This proves that $\mathcal{M}$ is $T^{*}_{z,B_{n}}$-invariant. ∎
The converse of the above observation is, of course, not true. For instance, consider $B_{n}$ as $B_{n}=zB_{n-1}$, where $B_{n-1}(0)=0$. Consider the $3$-dimensional subspace $\mathcal{M}$ of $H^{2}$, where
$\mathcal{M}=span\{B_{n-1},B_{n},zB_{n}\}.$
It is simple to see that $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant but neither nearly $T_{z}^{*}$-invariant nor $T_{z}^{*}$-invariant.
In terms of more examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces, we note that a class of Schmidt subspaces of Hankel operators are nearly $T_{z}^{*}$-invariant subspaces [6]. Given the above illustrations, one would naturally predict that the same would apply to finite-rank perturbations of Hankel operators or perturbations of operators that are related to Hankel operators. It would certainly be fascinating to identify such class operators.
Acknowledgement: The research of the first named author is supported by the Theoretical Statistics and Mathematics Unit, Indian Statistical Institute, Bangalore, India. The research of the second named author is supported in part by TARE (TAR/2022/000063) by SERB, Department of Science & Technology (DST), Government of India.